OpenStudy (anonymous):

Help! I must solve sin(3ø+π/18)=1 ø is theta

OpenStudy (anonymous):

to be more clear. I have set all values inside parenthesis to Alpha. its then properly redistributing those values that I am having trouble with..

OpenStudy (anonymous):

i am confused on where the 3 went

OpenStudy (anonymous):

Oh Sorry I did not see that \[ \sin(3\theta+ \frac \pi {18})=1=\sin(\frac \pi 2)\\ 3\theta+ \frac \pi {18}=\frac \pi 2 + 2 k \pi , \quad k=0, 1, 2, \cdots\\ 3\theta=\frac \pi 2 -\frac \pi {18}+ 2 k \pi , \quad k=0, 1, 2, \cdots\\ 3\theta=\frac {7\pi }{18}+ 2 k \pi , \quad k=0, 1, 2, \cdots\\ \theta=\frac {7\pi }{54}+ 2 k\frac \pi 3 , \quad k=0, 1, 2, \cdots\\ \]

OpenStudy (anonymous):

oh wow sweet. my teacher makes this so much more confusing than it needs to be. he thinks we're all idiots. he makes us set the value in the parenthesis to Alpha and then bring it back in at the end of the calculation.. its so annoying!

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

oh hey I forgot to add.. My bounded domain is -2π≤ø≤2π